# UNetMamba: An Efficient UNet-Like Mamba for Semantic Segmentation of High-Resolution Remote Sensing Images

### v2

@@ Abstract
Semantic segmentation of high-resolution remote sensing images is vital in downstream applications such as land-cover mapping, urban planning and disaster assessment. Existing Transformer-based methods suffer from the constraint between accuracy and efficiency, while the recently proposed Mamba is renowned for being efficient. Therefore, to overcome the dilemma, we propose UNetMamba, a UNet-like semantic segmentation model based on Mamba. It incorporates a mamba segmentation decoder (MSD) that can efficiently decode the complex information within high-resolution images, and a local supervision module (LSM), which is train-only but can significantly enhance the perception of local contents. Extensive experiments demonstrate that UNetMamba outperforms the state-of-the-art methods with mIoU increased by 0.87% on LoveDA and 0.36% on ISPRS Vaihingen, while achieving high efficiency through the lightweight design, less memory footprint and reduced computational cost. The source code is available at https://github.com/EnzeZhu2001/UNetMamba.
@@
As the foundation models of computer vision continuously evolve, research on semantic segmentation of high-resolution remote sensing images has flourished. The CNN-based method UNet[1] is a milestone in semantic segmentation, which has established the foundational architecture for subsequent works: an encoder-decoder framework with skip connections. With the success of Transformer[2], Wang et al.[3] introduce it into the U-shape framework, significantly improving the semantic segmentation accuracy of high-resolution remote sensing images. Consequently, numerous subsequent works have focused on innovating Transformer-based methods[4, 5, 6, 7]. Despite the prosperity of Transformer in remote sensing semantic segmentation, its quadratic computational complexity and high parameter count[8] have inevitably limited the application efficiency in high-resolution images.
Recently, Mamba [9] has become a hot point. For efficiency, it successfully achieves a linear complexity, while guaranteeing accuracy through its competitive long-distance dependency modeling capability. Subsequently, Mamba-based visual foundation models[10, 11] have been well-designed. Benefiting from such models, a series of Mamba-based UNets[12, 13, 14] have achieved impressive accuracy in medical imagery segmentation, showing the potential of visual Mamba in segmentation. For remote sensing image interpretation, the 2-D selective scanning of VMamba[10] has also demonstrated promising efficiency, with RSMamba[15] for classification, CDMamba[16] for change detection, and RS3Mamba[17] for semantic segmentation. However, these pioneering works have primarily validated the effectiveness of Mamba in their respective tasks, without fully leveraging its dual advantages in accuracy and efficiency through targeted design.
@@ II Methodology
As shown in Fig. 1(a), the overall architecture of UNetMamba is constructed under a U-shape framework [1], where the ResT backbone[18] is selected as encoder while decoder is the proposed MSD. Moreover, the LSM is designed to enhance the perception of local semantic information.
II-A Mamba Segmentation Decoder
In previous studies[12, 13, 14, 17], the benefits of employing VMamba[10] for remote sensing semantic segmentation have been established. However, models that utilize a pre-trained VMamba backbone while encoding tend to be overweight for our goal of high efficiency. Therefore, we incorporate the basic unit of VMamba, namely, VSS block[10], into the decoding side only.
The structure of a VSS block is depicted in Fig. 1(b). The 2-D feature map , which has undergone patch expansion, firstly goes through layer normalization and then splits into two information flows. The main flow is further linearly embedded as :
@@ II-A Mamba Segmentation Decoder
II-B Local Supervision Module
Due to the large receptive field of VSS blocks, MSD tends to partially overlook local semantic information while decoding . However, such details in high-resolution images are key factors to further improve the accuracy of semantic segmentation. Therefore, the LSM is proposed to enhance the perception of local details, as depicted in Fig. 1(c).
@@ II-B Local Supervision Module
BANet[3] | ResT-Lite | 12.73 | 194.61 | 85.43 | 90.30 | 82.43 | 91.93 |
ABCNet[4] | ResNet-18 | 13.67 | 206.74 | 62.62 | 90.36 | 82.49 | 91.94 |
MANet[5] | ResNet-50 | 35.86 | 547.87 | 216.82 | 90.65 | 83.04 | 92.26 |
DC-Swin[6] | Swin-T | 45.63 | 694.62 | 190.04 | 90.66 | 83.07 | 92.28 |
UNetFormer[7] | ResNet-18 | 11.72 | 179.34 | 47.03 | 90.56 | 82.89 | 92.20 |
RS3Mamba [17] | R18-VMamba-T | 43.32 | 662.21 | 157.89 | 90.71 | 83.06 | 92.23 |
UNetMamba(ours) | ResT-Lite | 14.76 | 225.71 | 100.52 | 90.91 | 83.43 | 92.46 |
@@
where denotes the kernel size, while denotes the output feature of MSD at the corresponding stage. The two branches are then merged by addition:
@@
During the training process, we add LSM blocks with different input channels to the decoder at stage 2-4 and then attain the segmentation result by summing up, which is further used in calculating the auxiliary loss function.
II-C Loss Functions
To ensure the holistic segmentation performance of UNetMamba, we select two classical loss functions to jointly form our principal loss:
where and represents dice loss and cross-entropy loss, respectively. And denotes the number of samples while denotes the number of categories. is the -th element in one-hot encoding of the true semantic labels for sample , while is the confidence of sample belonging to category . Moreover, to achieve the supervisory role of LSM, cross-entropy loss is also selected as the auxiliary loss for local supervision. Therefore, the overall loss function is formulated as follows:
@@ III-A Datasets
1) LoveDA: The LoveDA dataset[19] contains 5987 high-resolution remote sensing images at the same size of 1024 × 1024 pixels and includes 7 land-cover categories (background, building, road, water, barren, forest and agriculture). The dataset encompasses two scenes (urban and rural) collected from three Chinese cities (Nanjing, Changzhou and Wuhan), which brings considerable challenges due to multi-scale land-covers and complex contents. In our experiments, following the official settings[19], 2522 images were utilized for training, 1669 images for validation and 1796 images for testing without further cropping.
2) ISPRS Vaihingen: The ISPRS Vaihingen dataset contains 33 TOP (True OrthoPhoto) images, each with very high spatial resolution at an average size of 2494 × 2064 pixels. This widely-used remote sensing dataset involves 6 land-cover categories (impervious surface, low vegetation, building, tree, car and background). In the experiments, we randomly selected 9 images (area ID: 1, 5, 11, 15, 21, 26, 30, 33, 38) for testing, with the remaining 24 images for training, and cropped all of the images into patches at a size of 1024 × 1024 pixels.
@@ III-C Performance Comparison
We conducted a comprehensive performance comparison on LoveDA and ISPRS Vaihingen datasets. A series of SOTA optical remote sensing imagery semantic segmentation models were selected as our competitors, including Transformer-based models: BANet[3], ABCNet[4], MANet[5], DC-Swin[6], UNetFormer[7], and our baseline, RS3Mamba[17].
1) Comparison on the LoveDA Dataset: As shown in Tab. I, the proposed UNetMamba achieved the best accuracy performance in quantitative comparison, with a significant improvement of 0.87% in mIoU. It is note-worthy that our UNetMamba achieved SOTA in six out of seven categories. In background and agriculture, UNetMamba led the second place by 1.61% and 2.39%, respectively. This distinct advantage indicated that, benefiting from MSD, our UNetMamba garnered excellent global perception ability, which made it capable of accurately segmenting large land-covers under such a high-resolution condition. The outstanding performance of UNetMamba in building, road and other small land-cover categories proved that the train-only LSM successfully enhanced the perception of local contents through supervision and generalization. The qualitative comparison results on LoveDA are illustrated in Fig. 2, which intuitively shows that the semantic segmentation mask obtained by UNetMamba is more accurate with less omission and clearer boundaries.
2) Comparison on the ISPRS Vaihingen Dataset: The proposed UNetMamba not only achieved SOTA performance on the challenging LoveDA dataset, but also performed noticeably well on the classic Vaihingen dataset. As shown in Tab. II, in terms of accuracy, our UNetMamba achieved improvements of 0.20% in mF1, 0.36% in mIoU, and 0.18% in OA, while maintaining competitive efficiency through a lightweight model (14.76M) with low costs (225.71MB, 100.52G). Fig. 3 illustrates the qualitative comparison results, with the framed areas further demonstrating the competitiveness of UNetMamba in the perception of local semantic details, which is particularly important for the semantic segmentation accuracy in high-resolution remote sensing scenes.
@@
Baseline with ResT-Lite | 42.14 | 183.64 | 52.64 | 83.21 | |
✓ | 13.89 | 100.52 | 52.61 | 83.15 | |
✓ | ✓ | 14.76 | 100.52 | 53.35 | 83.43 |
@@ III-D Ablation Studies
To assess the effectiveness of the proposed MSD and LSM in UNetMamba, ablation studies were conducted on both of the datasets, with the results listed in Tab. III. For the sake of credibility, the baseline model, RS3Mamba[17], was updated by replacing its main encoding branch with the ResT[18] backbone consistent with UNetMamba. And after removing its Mamba-based auxiliary branch and upgrading its decoder into our MSD, the model experienced significant reduction in Param and FLOPs, with only minor drop in mIoUs, demonstrating the high efficiency of MSD in decoding complex semantic information. On both datasets, deployment of LSM also led to 0.74% and 0.28% increase in mIoUs, respectively. However, the cost was merely a modest parameter-count increase of 0.87M and a constant computation complexity due to the train-only scheme, which further confirming the high efficiency of LSM in enhancing the perception of local semantic information.

### v3

@@ Abstract
Semantic segmentation of high-resolution remote sensing images is vital in downstream applications such as land-cover mapping, urban planning and disaster assessment. Existing Transformer-based methods suffer from the constraint between accuracy and efficiency, while the recently proposed Mamba is renowned for being efficient. Therefore, to overcome the dilemma, we propose UNetMamba, a UNet-like semantic segmentation model based on Mamba. It incorporates a mamba segmentation decoder (MSD) that can efficiently decode the complex information within high-resolution images, and a local supervision module (LSM), which is train-only but can significantly enhance the perception of local contents. Extensive experiments demonstrate that UNetMamba outperforms the state-of-the-art methods with mIoU increased by 0.87% on LoveDA and 0.39% on ISPRS Vaihingen, while achieving high efficiency through the lightweight design, less memory footprint and reduced computational cost. The source code is available at https://github.com/EnzeZhu2001/UNetMamba.
@@
As the foundation models of computer vision continuously evolve, research on semantic segmentation of high-resolution remote sensing images has flourished. The CNN-based method UNet[1] is a milestone in semantic segmentation, which has established the foundational architecture for subsequent works: an encoder-decoder framework with skip connections. With the success of Transformer[2], Wang et al.[3] introduce it into the U-shape framework, significantly improving the semantic segmentation accuracy of high-resolution remote sensing images. Consequently, numerous subsequent works have focused on innovating Transformer-based methods[4, 5, 6]. Despite the prosperity of Transformer in remote sensing semantic segmentation, its quadratic computational complexity and high parameter count[7] have inevitably limited the application efficiency in high-resolution images.
Recently, Mamba [8] has been proposed and attracted extensive attention. For efficiency, it achieves a linear complexity, while guaranteeing accuracy through its competitive long-distance dependency modeling capability. Subsequently, Mamba-based visual foundation models[9, 10] have been well-designed. Benefiting from such models, a series of Mamba-based UNets[11, 12, 13] have achieved impressive accuracy in medical imagery segmentation, showing the potential of visual Mamba in segmentation. For remote sensing imagery interpretation, visual Mamba has also demonstrated promising efficiency, with RSMamba[14] for classification, CDMamba[15] for change detection, and RS3Mamba[18] for semantic segmentation. However, these pioneering works have primarily validated the effectiveness of Mamba in their respective tasks, without fully leveraging its dual advantages in accuracy and efficiency through targeted design.
@@ II Methodology
As shown in Fig. 1(a), the overall architecture of UNetMamba is constructed under a U-shape framework [1], where a pre-trained ResT backbone[19] is selected as encoder while decoder is the proposed MSD. Moreover, the LSM is designed to enhance the perception of local semantic information.
II-A ResT Encoder
As illustrated in Fig. 1(a), with Efficient Transformer Block (ETB) as the core, ResT encoder is consisted of four stages to capture multi-scale feature maps. To avoid the quadratic computational costs of vanilla multi-head self-attention, ETB employs efficient multi-head self-attention (EMSA)[19]:
where an instance normalization IN(·) and a 1×1 convolution Conv(·) are introduced to improve efficiency. When applied to high-resolution remote sensing images, EMSA can ensure the high encoding efficiency of our UNetMamba.
II-B Mamba Segmentation Decoder
Since high-resolution remote sensing images can be treated as very long sequences when flattened, the linear scaling ability of Mamba[8] has a natural advantage in processing them compared with the quadratic one of Transformer[2].
Actually, in pioneering studies[11, 12, 13, 18], the benefits of employing Mamba for semantic segmentation have been established. However, models that utilize a pre-trained VMamba[9] backbone (30M) while encoding tend to be overweight for our goal of high efficiency. Therefore, we incorporate the basic unit of VMamba, namely, VSS block[9], into the decoding side only.
The structure of a VSS block is depicted in Fig. 1(b). The 2-D feature map , which has undergone patch expansion, firstly goes through layer normalization and then splits into two flows. The main flow is further linearly embedded as :
@@ II-B Mamba Segmentation Decoder
| BANet[3] | ResT-Lite | 12.73 | 194.61 | 85.43 | 90.32 | 82.45 | 91.92 |
| MANet[4] | ResNet-50 | 35.86 | 547.87 | 216.82 | 90.68 | 83.06 | 92.28 |
| DC-Swin[5] | Swin-T | 45.63 | 694.62 | 190.04 | 90.71 | 83.08 | 92.30 |
| UNetFormer[7] | ResNet-18 | 11.72 | 179.34 | 47.03 | 90.59 | 82.93 | 92.21 |
| E-PyramidMamba[16] | ResNet-18 | 28.76 | 439.15 | 75.97 | 90.74 | 83.08 | 92.29 |
| CM-UNet[17] | ResNet-18 | 13.11 | 199.96 | 48.66 | 90.32 | 82.60 | 92.02 |
| RS3Mamba [18] | R18-VMamba-T | 43.32 | 662.21 | 157.89 | 90.73 | 83.08 | 92.27 |
| UNetMamba(ours) | ResT-Lite | 14.76 | 225.71 | 100.52 | 90.95 | 83.47 | 92.51 |
II-C Local Supervision Module
The large receptive field of SS2D in VSS blocks tend to be double-edged, causing MSD to partially overlook local semantic information while decoding. However, such details in high-resolution images are key factors to further improve the accuracy of semantic segmentation. Therefore, a CNN-based flexible LSM is proposed to enhance the perception of local details, as depicted in Fig. 1(c).
@@ II-C Local Supervision Module
where denotes the kernel size, while denotes the output feature of MSD at the corresponding stage. The two branches then merge into .
@@
During training, we add LSM blocks with different input channels to the decoder at stage 2-4 and then attain the segmentation result by summing up, which is further used in calculating the auxiliary loss function.
II-D Loss Function
The loss function utilized in training UNetMamba consists of two parts, a principal loss for overall optimization and an auxiliary loss for local supervision.
To guarantee the holistic segmentation performance of UNetMamba, two classic loss functions are selected to jointly form the principal loss :
where and represents dice loss and cross-entropy loss, respectively. And denotes the number of samples while denotes the number of categories. is the -th element in one-hot encoding of the true label for sample , while is the confidence of sample belonging to category .
Moreover, in order to efficiently achieve the local supervisory role of LSM, the most classic loss function in semantic segmentation [1] is also used as the auxiliary loss .
Eventually, to balance the effects of and and then achieve optimal performance, the overall loss function for training UNetMamba is formulated as a weighted equation:
@@ III-A Datasets
1) LoveDA: The LoveDA dataset[20] contains 5987 high-resolution remote sensing images at the same size of 1024×1024 pixels and includes 7 land-cover categories (background, building, road, water, barren, forest and agriculture). The dataset encompasses two scenes (urban and rural) collected from three Chinese cities, which brings considerable challenges due to multi-scale land-covers and complex contents. In our experiments, following the official settings[20], 2522 images were utilized for training, 1669 images for validation and 1796 images for testing without further cropping.
2) ISPRS Vaihingen: The ISPRS Vaihingen dataset contains 33 TOP (True OrthoPhoto) images, each with very high spatial resolution at an average size of 2494×2064 pixels. This widely-used remote sensing dataset involves 6 land-cover categories (impervious surface, tree, building, low vegetation, car and background). In the unbiased experiments, we randomly selected 9 images for testing, 2 images for validating with the remaining images for training, which was repeated for 5 times, and cropped all the 5 different sets of images into patches at a size of 1024×1024 pixels, respectively.
@@ III-C Performance Comparison
We conducted a comprehensive performance comparison on LoveDA and ISPRS Vaihingen datasets. A series of SOTA optical remote sensing imagery semantic segmentation models were selected as our competitors, including Transformer-based models: BANet[3], MANet[4], DC-Swin[5], UNetFormer[6], and Mamba-based models: Efficient PyramidMamba (E-PyramidMamba)[16], CM-UNet[17], RS3Mamba[18].
1) Comparison on the LoveDA Dataset: As shown in Tab. I, the proposed UNetMamba achieved the best accuracy performance in quantitative comparison, with a significant improvement of 0.87% in mIoU. It is note-worthy that our UNetMamba achieved SOTA in six out of seven categories. In background and agriculture, UNetMamba led the second place by 1.48% and 2.39%, respectively. This distinct advantage indicated that, benefiting from MSD, our UNetMamba garnered excellent global perception ability, which made it capable of accurately segmenting large land-covers under such a high-resolution condition. The outstanding performance of UNetMamba in building, road and other small land-cover categories proved that the train-only LSM successfully enhanced the perception of local contents through supervision and generalization. The qualitative comparison results on LoveDA are illustrated in Fig. 2, which intuitively shows that the semantic segmentation mask obtained by UNetMamba is more accurate with less omission and clearer boundaries.
2) Comparison on the ISPRS Vaihingen Dataset: The proposed UNetMamba not only achieved SOTA performance on the challenging LoveDA dataset, but also performed noticeably well on the classic Vaihingen dataset. As shown in Tab. II, in terms of accuracy, our UNetMamba achieved improvements of 0.21% in mF1, 0.39% in mIoU, and 0.21% in OA, while maintaining competitive efficiency through a lightweight model (14.76M) with low costs (225.71MB, 100.52G). Fig. 3 illustrates the qualitative comparison results, with the framed areas further demonstrating the competitiveness of UNetMamba in the perception of local semantic details, which is particularly important for the semantic segmentation accuracy in high-resolution remote sensing scenes.
@@
| Baseline with ResT-Lite | 42.14 | 183.64 | 52.64 | 83.23 | |
| ✓ | 13.89 | 100.52 | 52.61 | 83.18 | |
| ✓ | ✓ | 14.76 | 100.52 | 53.35 | 83.47 |
@@ III-D Ablation Studies
To assess the effectiveness of the proposed MSD and LSM in UNetMamba, ablation studies were conducted on both of the datasets, with the results listed in Tab. III. For the sake of credibility, the baseline model, RS3Mamba[17], was updated by replacing its main encoding branch with the ResT[18] backbone consistent with UNetMamba. And after removing its Mamba-based auxiliary branch and upgrading its decoder into our MSD, the model experienced significant reduction in Param and FLOPs, with only minor drop in mIoUs, demonstrating the high efficiency of MSD in decoding complex semantic information. On both datasets, deployment of LSM also led to 0.74% and 0.29% increase in mIoUs, respectively. However, the cost was merely a modest parameter-count increase of 0.87M and a constant computation complexity due to the train-only scheme, which further confirming the high efficiency of LSM in enhancing the perception of local semantic information.